# Empowering LLMs to Understand and Generate Complex Vector Graphics

### v2

@@ Abstract
The unprecedented advancements in Large Language Models (LLMs) have profoundly impacted natural language processing but have yet to fully embrace the realm of scalable vector graphics (SVG) generation. While LLMs encode partial knowledge of SVG data from web pages during training, recent findings suggest that semantically ambiguous and tokenized representations within LLMs may result in hallucinations in vector primitive predictions. Additionally, LLM training typically lacks modeling and understanding of the rendering sequence of vector paths, which can lead to occlusion between output vector primitives. In this paper, we present LLM4SVG, an initial yet substantial step toward bridging this gap by enabling LLMs to better understand and generate vector graphics. LLM4SVG facilitates a deeper understanding of SVG components through learnable semantic tokens, which precisely encode these tokens and their corresponding properties to generate semantically aligned SVG outputs. Using a series of learnable semantic tokens, a structured dataset for instruction following is developed to support comprehension and generation across two primary tasks. Our method introduces a modular architecture to existing large language models, integrating semantic tags, vector instruction encoders, fine-tuned commands, and powerful LLMs to tightly combine geometric, appearance, and language information. To overcome the scarcity of SVG-text instruction data, we developed an automated data generation pipeline that collected a massive dataset of more than 250k SVG data and 580k SVG-text instructions, which facilitated the adoption of the two-stage training strategy popular in LLM development. By exploring various training strategies, we developed LLM4SVG, which significantly moves beyond optimized rendering-based approaches and language-model-based baselines to achieve remarkable results in human evaluation tasks. Code, model, and data will be released at: https://ximinng.github.io/LLM4SVGProject/
@@ 1 Introduction
Scalable Vector Graphics (SVGs) represent a fundamental method of image encoding, where visual elements are formed from basic shapes like vector paths, curves, and polygons, defined by mathematical equations. This contrasts with raster graphics, which depict images as pixel arrays on a grid. One of the key advantages of vector graphics is their capacity to preserve high precision and consistent visual quality across varying resolutions, as they can be scaled indefinitely without any loss of detail. Additionally, vector graphics typically maintain compact file sizes, which enhances their efficiency for storage and transmission. More critically, they offer superior editability, enabling designers to easily select, modify, and compose elements. This attribute is particularly vital during the design process, facilitating seamless adjustments and fostering creative exploration. Due to their inherent properties, SVGs are particularly well-suited for applications in visually fluent design.
@@
Recent advancements in large language models (LLMs) [39, 34, 2, 55, 28] have evidenced their capacity to comprehend and parse XML syntax [5], achieved through extensive pre-training on a diverse corpus of text data sourced from the Web. This proficiency establishes a robust foundation for LLMs to synthesize vector graphics [43]. Notably, state-of-the-art models, such as ChatGPT [34] and GPT-4 [2], show proficiency in generating simple vector primitives like triangles and rectangles; however, they often encounter significant limitations in synthesizing complex graphics. This is because, during pre-training, SVG data from the Internet is often embedded within web page code, requiring LLMs to parse layered languages, which renders SVG data less accessible amidst lengthy XML tags. These limitations manifest as confusion or hallucinations in vector path sequences, yielding semantically ambiguous graphics and improperly encoded attributes.
In this paper, we present LLM4SVG, an initial yet substantial step toward bridging this gap by enabling LLMs to better understand and generate vector graphics, as illustrated in Fig. 1. Built on an existing LLM, such as GPT-2 [39] and LLaVA [28], LLM4SVG maximizes the model’s potential for vector graphic synthesis. Our paper makes the following contributions:
  • Introducing LLM4SVG: Advancing LLM Capabilities for SVG Understanding and Generation. We present LLM4SVG, a pioneering model designed to bridge the gap in SVG generation within large language models (LLMs), addressing challenges in accurately predicting vector primitives and managing occlusions. By utilizing learnable semantic tokens, LLM4SVG enables LLMs to effectively encode SVG components and properties, aligning generation outputs with human design standards and reducing the risk of semantic ambiguities.
  • Modular Architecture Integrating Geometric, Visual, and Linguistic Features. LLM4SVG features a modular design that augments traditional LLM architectures with a vector instruction encoder, semantic token embedding, and regression heads for precise coordinate and color predictions. This multimodal integration facilitates a comprehensive understanding of SVG elements, enabling LLMs to generate refined, semantically consistent vector graphics by effectively merging geometry, appearance, and linguistic information.
  • Development of a Large-Scale SVG-Text Dataset and a Two-Stage Training Strategy. Addressing the lack of SVG-Text instruction data, LLM4SVG introduces an automated pipeline that collects 580k SVG-text instruction pairs. We collected 250k SVG samples designed by human designers. This large-scale dataset supports a two-stage training strategy, which enhances LLM comprehension and performance in SVG generation tasks, setting a robust foundation for further advancements in SVG-related capabilities within LLMs.
2 Related Word
2.1 Vector Graphics Synthesis
@@ 2.2 Large Language Models
Large Language Models (LLMs) have achieved substantial progress in natural language understanding, exhibiting strong generalization and reasoning capabilities to complete real-world applications through extensive pre-training on large-scale textual datasets, such as ChatGPT [34]/GPT-4 [2], LLaMA [55], Claude 3.5 [35], Qwen 2.5 [52], Gemini [7], Grok-2 [65], Yi [70], Falcon [3], Phi series [32, 24, 1], etc.
Broadly, LLMs can be classified into two categories. The first category includes models that employ a large language model to interface with individual, modality-specific models [15, 21, 37]. This approach circumvents the need for model training but is heavily dependent on the availability and capabilities of pre-existing models or APIs. The second category includes models that use an end-to-end training strategy, comprising two main paradigms. The first involves training the model from scratch, similar to text-only LLMs, using large-scale multi-modal corpora and datasets [22, 2]. Instead of training an LLM from scratch, it is more efficient and practical to start with a pre-trained one [28]. In our work, we adopt a fine-tuning strategy based on pre-trained LLMs, focusing on adapting the model to generate and understand vector graphics. While multi-modal capabilities can be incorporated when needed, our primary emphasis is on leveraging the pre-trained LLM’s robust foundational knowledge and generalization abilities, which provide a strong basis for specialized, task-oriented applications in the vector graphics domain.
The adaptability of LLMs also extends naturally into the multi-modal domain. This has enabled them to be utilized not only for text-based tasks but also as foundational architectures for expanding into multi-modal applications. Building on this foundation, Multi-modal Large Language Models (MLLMs) are designed to interpret and generate content across various modalities, including but not limited to images [13, 22, 28, 60, 74, 59], audio [8, 21], motion [25, 72], and 3D point clouds [18, 68].
@@ 2.3 Instruction Tuning
In the field of natural language processing (NLP), researchers have investigated various methods [36, 62, 34] for instruction-tuning LLMs, aiming to enhance their ability to follow natural language instructions and perform real-world tasks. It has been found that this simple and straightforward approach can significantly enhance the zero-shot and few-shot generalization capabilities of LLMs. With the growing popularity of multimodal large language models (MLLMs), LLaVA [28] has utilized visual instruction tuning with vision-language instruction data, significantly enhancing the performance of open-source MLLMs on multimodal tasks.
Recently, several studies have explored directly fine-tuning LLMs with image embeddings for generating Scalable Vector Graphics (SVG) by treating SVG code as a text-based representation [43, 69]. While these methods have shown promise in applying LLMs to vector graphics generation, they often overlook the hierarchical and structured nature of SVG files, treating them as sequential text. Additionally, most LLMs [59, 52, 1] often treat numbers as individual characters, making it difficult for them to understand continuous numerical data, such as coordinates. This simplification can restrict the model’s ability to fully capture the SVG’s complex dependencies and file structure, which can impact the quality and semantic accuracy of the generated graphics.
3 Vector Instruction Following Data
@@
USER Generate an SVG based on the following text: $\{\mathrm{prompt}\}$ </s>
@@
# 2 SYSTEM You are a helpful assistant, please help me generate SVG by referring to this image </s>
USER Refer to this image: $\{\mathrm{img}\}$ and generate SVG based on the following text: $\{\mathrm{prompt}\}$ </s>
@@
# 4 SYSTEM Attempt to understand the path composition in this svg </s>
@@
# 5 SYSTEM Attempt to understand the path composition in this svg </s>
@@
SVG Re-captioning. SVG data collected from the Internet often contains noise, and directly using it for learning can compromise the model’s potential representational accuracy. Approximately half of the data in an SVG file is redundant for visual rendering. This redundancy includes: (1) temporary data used by vector editing applications, (2) non-optimal structural representations of SVG, and (3) unused and invisible graphic elements. We propose an SVG data preprocessing pipeline designed to losslessly reduce the size of SVG files generated by vector editing applications. Details are provided in Sec. B and Fig. S5 of Supplementary. This pipeline incorporates a combination of carefully optimized strategies that address four key aspects of SVG files: elements, attributes, paths, and outputs. After optimizing the SVG, we rasterize it into an image of $512\times 512$ pixels and use the BLIP [26] model to generate a corresponding caption as a text prompt. Consequently, we obtain a multimodal dataset, each entry of which is a triplet consisting of the optimized SVG, its corresponding rasterized image, and a text description generated by the BLIP model.
@@
SVG Instruction Data. The constructed dataset adheres to a standardized instruction format, as depicted in Table 1 #1, comprising Text-SVG pairs for fine-tuning in text-to-SVG generation tasks. For text-guided SVG synthesis, visual prompts are indispensable. As illustrated in Table 1 #2, Text-Image-SVG triples facilitate instruction tuning for text-and-image to SVG generation.
During the SVG re-captioning phase, we obtained the corresponding text descriptions from the BLIP based on the rendering results, which were sufficient for text prompts, but too short for comments to understand SVG. Inspired by the recent success of GPT models in text annotation tasks [12], we utilized ChatGPT/GPT-4 for instruction-following data generation.
@@ 4.1 SVG Semantic Tokens
For an input SVG $\mathbf{X}_{v}$, we convert it from raw code into a structured representation. To accomplish this, we defined 55 SVG semantic tokens $s_{i}\in\mathbb{R}^{55}$ (including 15 tag tokens, 30 attribute tokens, 10 path command tokens, details are provided in Sec. D and Fig.  S1 of Supplementary). These SVG tokens are used to replace all tags and attributes in the SVG source code, thus preventing the textual encoding of SVG tags and attributes as regular text. For example, the tag <path> will be tokenized as an SVG semantic token, rather than the literal “path” by the tokenizer. This ensures the uniqueness of SVG tags and attributes, and allows for their efficient integration into LLMs in a manner that is consistent with SVG definitions and optimizes token initialization. We adapt the embedding layer $\mathbf{W}_{\mathrm{emb}}\in\mathbb{R}^{|\mathcal{V}|^{\prime}}$ to learn the embeddings of these new tokens, where $|\mathcal{V}|^{\prime}:=|\mathcal{V}|+55$ represents the sum of the size of the original vocabulary and the additional SVG tokens. Each new token is initialized based on the semantic average of its descriptive text, as defined by the equation:
$\bm{E}(s_{i})=\frac{1}{n}\sum_{j=1}^{n}\mathbf{W}_{\mathrm{emb}}^{\top}\cdot w_{i,j}$ (1)
where $w_{i,j}$ represents the $j$-th description token of $i$-th $s_{i}$, $\bm{E}(\cdot)$ represents the token embedding layer and $n=55$. This initialization provides a good starting point for each SVG token and builds a compact, distributed representation for all SVG tokens.
@@ 4.2 Architecture
[…]
@@
4.3 Training
Stage 1: Pre-training for Feature Alignment. These pairs are converted to instruction-following data using the naive expansion method describe in Sec. 3. Each sample can be treated as a single-turn conversation. To construct the input $\mathbf{X}_{\mathrm{inst}}$, instruction data is randomly sampled. In training, we keep the weights of both the visual encoder and the LLM frozen, and focus on maximizing the likelihood of Eq. 2 with trainable parameters $\theta=\mathbf{W}_{\mathrm{emb}}$ (the word embedding) only.
5 Experiments
Overview. This section outlines the core aspects of our model implementation, followed by a description of the dataset collection and preprocessing pipeline, which involves 250,000 high-quality SVG files. We then empirically evaluate the effectiveness of our model, LLM4SVG, in generating high-quality SVGs. This evaluation benchmarks our model against current state-of-the-art methods both qualitatively and quantitatively, as detailed in Sec. 5.1 and  5.2. This evaluation is augmented with a comprehensive architectural analysis, including ablation studies detailed in Sec. 5.3, to identify the specific contributions of individual model components to the overall performance.
Figure 3: Qualitative Comparison Between LLM4SVG and State-of-the-Art SVG Generation Methods, including optimization-based and LLM-based methods.
Implementation Details. In the initial stage of training, we exclusively focus on the word embeddings $\mathbf{W}_{\mathrm{emb}}$. As mentioned in Sec. 4.1, we added 55 SVG semantic tokens and initialized the word embedding using the semantic average of the description text.
In the second stage, we employ an end-to-end fine-tuning strategy to comprehensively optimize model performance. For the LoRA [19] and QLoRA [9] approaches, we configure the LoRA rank $r=32$ and scaling factor $\alpha=32$.
During both training stages, we apply the AdamW optimizer with hyper-parameters $\beta_{1}=0.9,\beta_{2}=0.999,\epsilon=1\times 10^{-8}$, and a learning rate of $lr=2\times 10^{-5}$. We also address the potential complexity of certain SVG structures by setting a maximum token length of 2048. If an SVG’s token sequence exceeds this length, it is directly truncated to the maximum length. This approach prevents excessive sequence lengths that could hinder training efficiency and ensures the model remains within manageable computational limits. Training was performed using 8 NVIDIA A800 GPUs.
@@ 5 Experiments
Datasets. Our dataset comprise 250,000 emoji and icon SVGs, sourced from TwEmoji [56], NotoEmoji [14], FluentUIEmoji [31], as well as SVGRepo [50] and Reshot [42] websites. Specifically, the TwEmoji, NotoEmoji, and FluentUIEmoji subsets each contain approximately 3,000–4,000 SVGs. The Reshot website contributes 30,000 SVGs, while SVGRepo provides the remaining 210,000. As mentioned earlier, after collecting the SVG files, we preprocess them to enhance training efficiency and ensure consistency. Evaluation Metrics. To facilitate a comprehensive assessment of our proposed method compared to baseline approaches, we classify all methods into three categories: Optimization-based methods, Neural Network-based methods, and LLM-based methods. We then evaluate these methods across two key dimensions: visual quality and computational cost. For visual quality, we measure (1) visual quality using FID (Frechet Inception Distance)[17]; (2) text prompt alignment through the CLIP score [40]; and (3) aesthetic appeal using the Aesthetic score [46] and HPS (Human Preference Scores) [64]. (4) for computational cost, we test and compare the time cost of generating 10 SVGs by each method. (5) Avg.Tok represents the length of SVG code after removing comments and whitespace.

### v3

@@ Abstract
The unprecedented advancements in Large Language Models (LLMs) have profoundly impacted natural language processing but have yet to fully embrace the realm of scalable vector graphics (SVG) generation. While LLMs encode partial knowledge of SVG data from web pages during training, recent findings suggest that semantically ambiguous and tokenized representations within LLMs may result in hallucinations in vector primitive predictions. Additionally, LLM training typically lacks modeling and understanding of the rendering sequence of vector paths, which can lead to occlusion between output vector primitives. In this paper, we present LLM4SVG, an initial yet substantial step toward bridging this gap by enabling LLMs to better understand and generate vector graphics. LLM4SVG facilitates a deeper understanding of SVG components through learnable semantic tokens, which precisely encode these tokens and their corresponding properties to generate semantically aligned SVG outputs. Using a series of learnable semantic tokens, a structured dataset for instruction following is developed to support comprehension and generation across two primary tasks. Our method introduces a modular architecture to existing large language models, integrating semantic tags, vector instruction encoders, fine-tuned commands, and powerful LLMs to tightly combine geometric, appearance, and language information. To overcome the scarcity of SVG-text instruction data, we developed an automated data generation pipeline that collected our SVGX-SFT Dataset, consisting of high-quality human-designed SVGs and 580k SVG instruction following data specifically crafted for LLM training, which facilitated the adoption of the supervised fine-tuning strategy popular in LLM development. By exploring various training strategies, we developed LLM4SVG, which significantly moves beyond optimized rendering-based approaches and language-model-based baselines to achieve remarkable results in human evaluation tasks. Code, model, and data will be released at: https://ximinng.github.io/LLM4SVGProject/
@@ 1 Introduction
Scalable Vector Graphics (SVGs) constitute a fundamental image encoding paradigm wherein visual elements are constructed from primitive geometric entities defined by mathematical formulations, contrasting with raster graphics’ discrete pixel matrices. Vector-based representation offers resolution independence, preserving geometric precision across arbitrary scaling transformations without degradation. Vector graphics exhibit superior compression efficiency, optimizing storage requirements and transmission bandwidth. Their parametric editability enables precise manipulation of constituent elements—a characteristic instrumental during iterative design processes. These mathematical underpinnings render SVGs exceptionally suitable for applications demanding visual fluency and precision.
@@
Recent advancements in large language models (LLMs) [36, 2, 4, 54, 31] have evidenced their capacity to comprehend and parse XML syntax [6], achieved through extensive pre-training on a diverse corpus of text data sourced from the Web. This proficiency establishes a robust foundation for LLMs to synthesize vector graphics [43]. Notably, state-of-the-art models, such as GPT-4 [2] and Claude [4], show proficiency in generating simple vector primitives like triangles and rectangles; however, they often encounter significant limitations in synthesizing complex graphics. This is because, during pre-training, SVG data from the internet is often embedded within web page code, requiring LLMs to parse layered languages, which renders SVG data less accessible amidst lengthy XML tags. These limitations manifest as confusion or hallucinations in vector path sequences, yielding semantically ambiguous graphics and improperly encoded attributes.
In this paper, we present LLM4SVG, an initial yet substantial step toward bridging this gap by enabling LLMs to better understand and generate vector graphics, as illustrated in Fig. 1. Built on an existing LLM/MLLM, LLM4SVG maximizes the model’s potential for vector graphic synthesis. Our paper makes the following contributions:
  • LLM4SVG: The First Framework Supporting Arbitrary LLMs for SVG Tasks. We present LLM4SVG, a framework enabling any LLM or MLLM to understand and generate SVGs effectively. Our approach addresses vector primitive prediction challenges through learnable semantic tokens, precisely encoding SVG components and properties. This bridges natural language processing and vector graphics, aligning outputs with human design principles.
  • Modular Architecture with Decoupled Vector Instructions and Parameters. LLM4SVG enhances traditional LLM architectures by decoupling vector instructions from parameters. This integration enables comprehensive understanding of SVG elements, producing semantically consistent vector graphics by merging geometry, appearance, and linguistic information.
  • Development of SVGX-SFT Dataset. We introduce the SVGX-SFT Dataset with high-quality human-designed SVGs and 580k instruction following data crafted for LLM training. This dataset supports our training strategy, enhancing model performance in SVG generation and establishing a foundation for future vector graphics research.
2 Related Work
2.1 Vector Graphics Generation
@@
2.2 Vector Graphics Understanding
Recent research in vector graphics has advanced both recognition and evaluation methodologies. YOLAT [28] pioneered treating vector graphics recognition as a detection problem without rasterization, though it struggles with complex hierarchical structures and semantic relationships. YOLAT++ [12] addressed these limitations by introducing hierarchical recognition capabilities and a new chart-based dataset. Meanwhile, VGBench [74] developed a comprehensive benchmark evaluating LLMs on vector graphics understanding and generation across various formats, revealing that while LLMs show promise, they perform poorly on low-level formats like SVG. Despite these advances, existing approaches remain constrained—recognition models lack generation capabilities, evaluation frameworks don’t address fundamental representation issues, and none adequately capture the bidirectional relationship between natural language and vector graphics necessary for creative design workflows.
@@ 2.3 Large Language Models
Large Language Models (LLMs) have made significant strides in natural language understanding, demonstrating strong generalization and reasoning abilities through extensive pre-training on large-scale text corpora [36, 2, 54, 4, 51, 9, 63, 68, 3, 34, 26, 1].
LLMs can be broadly categorized into two types. The first type serves as an interface for individual modality-specific models [17, 23, 38], eliminating the need for retraining but relying heavily on external model availability. The second type employs an end-to-end training approach, which can either train models from scratch using large-scale multi-modal datasets [24, 2] or fine-tune pre-trained LLMs for specific applications [31]. Our work follows the latter strategy, adapting pre-trained LLMs to generate and understand vector graphics while maintaining flexibility for multi-modal extensions.
LLMs naturally extend into multi-modal domains, enabling them to process and generate content beyond text. Multi-modal Large Language Models (MLLMs) leverage this capability to handle diverse modalities, including images [16, 24, 31, 58, 73, 57], audio [10, 23], motion [27, 70], and 3D point clouds [20, 66]. This adaptability makes LLMs a powerful foundation for expanding into vector graphics and other specialized tasks.
@@ 2.4 Instruction Tuning
In natural language processing, researchers have explored various methods [37, 60, 36] for instruction-tuning LLMs to improve their ability to follow natural language instructions and perform real-world tasks. This straightforward approach has been shown to significantly enhance LLMs’ zero-shot and few-shot generalization capabilities. With the rise of multimodal large language models, LLaVA [31] has leveraged visual instruction tuning with vision-language data, greatly improving open-source MLLMs’ performance on multimodal tasks.
Recently, studies have investigated fine-tuning LLMs with image embeddings to generate Scalable Vector Graphics (SVG) by treating SVG code as a text-based representation [43, 67]. While promising, these approaches often overlook the hierarchical and structured nature of SVG files, treating them merely as sequential text.
3 SVGX-SFT Dataset
@@
USER Generate an SVG illustration from the given description: $\{\mathrm{prompt}\}$ </s>
@@
# 2 SYSTEM You are a helpful assistant, please help me generate an SVG from this image and description. </s>
USER Refer to rendering image: $\{\mathrm{img}\}$ and generate SVG from the given description: $\{\mathrm{prompt}\}$ </s>
@@
# 4 SYSTEM Describe this SVG based on its image representation </s>
@@
# 5 SYSTEM Describe this SVG based on its image representation </s>
@@
Figure 2: Overview of SVGX-SFT Dataset. (a) Top Frequent Words in SVGX-SFT Dataset. The most frequently occurring words in the SVGX-SFT dataset, highlighting common patterns and terminology used in SVG metadata. (b) Comparison of SVG Element Counts Before and After Cleaning. A comparison of the number of SVG elements before and after the cleaning process. The reduction in element count demonstrates the effectiveness of our preprocessing steps. (c) Word Cloud of SVG Names and Descriptions. A word cloud visualization of SVG names and descriptions, illustrating the distribution and emphasis of different terms in the dataset. (d) Group (<g></g>) Number Distribution. The distribution of <g> (group) elements in the dataset, showing the frequency of grouped elements and their structural significance within SVG files.
@@
SVG Re-captioning. SVG data collected from the Internet often contains noise, and directly using it for learning can compromise the model’s potential representational accuracy. Approximately half of the data in an SVG file is redundant for visual rendering. This redundancy includes: (1) temporary data used by vector editing applications, (2) non-optimal structural representations of SVG, and (3) unused and invisible graphic elements. We propose an SVG data preprocessing pipeline designed to losslessly reduce the size of SVG files generated by vector editing applications. Details are provided in Sec. B and Fig. S5 of Supplementary. After optimizing the SVG, we rasterize it into an image of $512\times 512$ pixels and use the BLIP [29] model to generate a corresponding caption as a text prompt. Consequently, we obtain a multimodal dataset, each entry of which is a triplet consisting of the optimized SVG, its corresponding rasterized image, and a text description generated by the BLIP model.
@@
SVG Instruction Following Data. The constructed dataset adheres to a standardized instruction format, as depicted in Table 1 #1, comprising Text-SVG pairs for fine-tuning in text-to-SVG generation tasks. For text-guided SVG synthesis, visual prompts are indispensable. As illustrated in Table 1 #2, Text-Image-SVG triples facilitate instruction tuning for text-and-image to SVG generation.
During the SVG re-captioning phase, we obtained the corresponding text descriptions from the BLIP based on the rendering results, which were sufficient for text prompts, but too short for comments to understand SVG. Inspired by the recent success of GPT models in text annotation tasks [15], we utilized ChatGPT [36]/GPT-4 [2] for instruction-following data generation.
@@ 4.1 SVG Semantic Tokens
For an input SVG $\mathbf{X}_{v}$, we convert it from raw code into a structured representation. To accomplish this, we defined 55 SVG semantic tokens $s_{i}\in\mathbb{R}^{55}$ (including 15 tag tokens, 30 attribute tokens, 10 path command tokens, details are provided in Sec. C and Fig. S1 of Supplementary). These SVG tokens are used to replace all tags and attributes in the SVG source code, thus preventing the textual encoding of SVG tags and attributes as regular text. For example, the tag <path> will be tokenized as an SVG semantic token, rather than the literal “path” by the tokenizer. This ensures the uniqueness of SVG tags and attributes, and allows for their efficient integration into LLMs in a manner that is consistent with SVG definitions and optimizes token initialization. We adapt the embedding layer $\mathbf{W}_{\mathrm{emb}}\in\mathbb{R}^{|\mathcal{V}|^{\prime}}$ to learn the embeddings of these new tokens, where $|\mathcal{V}|^{\prime}:=|\mathcal{V}|+55$ represents the sum of the size of the original vocabulary and the additional SVG tokens. Each new token is initialized based on the semantic average of its descriptive text $s$, as defined by the equation:
$\bm{E}(s)=\frac{1}{n}\sum_{j=1}^{n}\mathbf{W}_{\mathrm{emb}}^{\top}\cdot w_{j}$ (1)
where $w_{j}$ represents the $j$-th description token of $s$, $\bm{E}(\cdot)$ represents the token embedding layer and $n$ represents the length of the token after encoding the description $s$. This initialization provides a good starting point for each SVG token and builds a compact, distributed representation for all SVG tokens.
@@ 4.2 Architecture
4.3 Training
Stage 1: Pre-training for Feature Alignment. These pairs are converted to instruction-following data using the naive expansion method describe in Sec. 3. Each sample can be treated as a single-turn conversation. To construct the input $\mathbf{X}_{\mathrm{inst}}$, instruction data is randomly sampled. In training, we keep the weights of both the visual encoder and the LLM frozen, and focus on maximizing the likelihood of Eq. 2 with trainable parameters $\theta=\mathbf{W}_{\mathrm{emb}}$ only, where $\mathbf{W}_{\mathrm{emb}}$ represents the word embedding.
5 Experiments
Overview. This section outlines the core aspects of our model implementation, followed by a description of the dataset collection and preprocessing pipeline, which involves 250,000 high-quality SVG files, as detailed in Sec. 5.1. We then empirically evaluate the effectiveness of our model, LLM4SVG, in generating high-quality SVGs. This evaluation benchmarks our model against current state-of-the-art methods both qualitatively and quantitatively, as detailed in Sec. 5.2 and  5.3. This evaluation is augmented with a comprehensive architectural analysis, including ablation studies detailed in Sec. 5.4, to identify the specific contributions of individual model components to the overall performance.
Figure 4: Qualitative comparison of LLM4SVG with state-of-the-art SVG generation methods, categorized into optimization-based and LLM-based approaches. The prompts (left) guide the generation of vector graphics across different methods. Optimization-based methods (left section) focus on refining SVGs through iterative optimization, while LLM-based methods (right section) generate SVGs directly from text descriptions. The results highlight differences in abstraction, structure, and fidelity to the input prompt, with our LLM4SVG achieving more structured and visually coherent outputs.
@@ 5 Experiments
(a) Quantitative Comparison between LLM4SVG and State-of-the-Art Text-to-SVG Methods.
@@
Implementation Details. Our training process consists of two steps. In the first step, we add 55 SVG semantic tokens (Tab. S1) to the tokenizer and initialize the word embeddings $\mathbf{W}_{\mathrm{emb}}$ using the semantic average of the description text as described in Eq. 1. In the second step, we perform supervised fine-tuning (SFT) using two alternative methods: either LoRA/QLoRA [21, 11] fine-tuning or full-parameter fine-tuning. The trainable parameters include the word embeddings and the parameters involved in SFT training. SFT training typically requires 1 to 3 epochs.
During the training process, we apply the AdamW optimizer with hyper-parameters $\beta_{1}=0.9,\beta_{2}=0.999,\epsilon=1\times 10^{-8}$. We use a learning rate of $lr=3\times 10^{-4}$ with a cosine scheduler type and warmup ratio of 0.1. The training runs for 2 epochs with a maximum token length of 4096. If an SVG’s token sequence exceeds this length, it is directly truncated to the maximum length to ensure computational efficiency. Our implementation is based on LlamaFactory framework [72]. Considering VRAM limitations, we also integrated Unsloth [8] to support efficient training of quantized models. The training was performed using 8 NVIDIA A800 GPUs.
@@
Evaluation Metrics. To facilitate a comprehensive assessment of our proposed method compared to baseline approaches, we classify all methods into three categories: Optimization-based methods, Neural Network-based methods, and LLM-based methods. We then evaluate these methods across two key dimensions: visual quality and computational cost. For visual quality, we measure (1) visual quality using FID (Fréchet Inception Distance) [19]; (2) text prompt alignment through the CLIP score [41]; and (3) aesthetic appeal using the Aesthetic score [46] and HPS (Human Preference Scores) [62]. (4) for computational cost, we test and compare the time cost of generating 10 SVGs by each method. (5) Avg.Tok represents the length of SVG code after removing comments and whitespace.
5.1 SVGX-SFT Dataset Analysis
Figure 2 provides a comprehensive breakdown of the SVGX-SFT dataset, which is designed for LLM4SVG through supervised fine-tuning. Fig. 2(a): The most common words in the dataset include “logo”, “document”, “button”, and “download”, suggesting that the dataset primarily consists of UI-related SVG elements. This indicates a strong presence of design-related components, making the dataset particularly useful for training models that need to understand or generate UI elements. Fig. 2(b): The dataset underwent a cleaning process, significantly reducing the number of <path> elements and path coordinates. This removes redundant or unnecessary elements, improving data quality and reducing complexity. However, certain elements like <rect>, <circle>, and <ellipse> remained relatively stable, indicating their importance in the dataset. Fig. 2(c): The word cloud visualization highlights key terms associated with SVG elements, reinforcing the dominant themes from subfigure (a). Words like “button”, “cloud”, “folder”, and “icon” emphasize the dataset’s focus on UI-related graphics. The presence of words such as “geometric”, “outline”, and “pattern” suggests that the dataset also includes a variety of structured vector graphics. Fig. 2(d): The pie chart shows the distribution of SVG group counts, with a significant portion (23.1%) having 5 groups, followed by other group counts such as 4, 6, and 3. This indicates that a considerable number of SVGs in the dataset are structured hierarchically, which is crucial for understanding compositional relationships in vector graphics.